In a Harris​ poll, adults were asked if they are in favor of abolishing the penny. Among the​ responses, 1210 answered​ "no," 458 answered​ "yes," and 360 had no opinion. What is the sample proportion of yes​ responses, and what notation is used to represent​ it?
Answer #1

Number that answered yes = 458
Total responses = (1210+458+360) = 2028

Proportion of yes responses = 458/2028≈ 0.2258 or 22.58%

Answer:
22.58% said yes; it's represented as a percentage.